OpenStudy (anonymous):

please helpppp!!!!!!!!!!! - Zach borrowed 4,000 from his grandfather to buy a used car so he could travel to and from work. After working for 2 years, he has saved 5,200 to pay back to his grandfather. At what interest rate did Zach borrow the money? Round to the nearest whole percent.

OpenStudy (anonymous):

5200/4000=x/100 520000=4000x 130=x not sure on that thugh